# Search for continuous gravitational waves from known pulsars in the first part of the fourth LIGO-Virgo-KAGRA observing run

### v1

@@ 4.2 EM dataset
For this analysis, we selected pulsars with a rotation frequency close to or greater than , to lie in the bandwidth of the LIGO detectors, and with an expected targeted search sensitivity for the strain amplitude within a factor 3 of the spin-down limit (see Figure 1 for the pulsars’ sky locations and Table Acknowledgement for the list of analyzed pulsars). Out of the 45 pulsars analyzed in this work, 11 pulsars belong to binary systems and there are 10 millisecond pulsars with frequencies higher than .
@@ 5.1 Targeted searches
The results are shown in Table Acknowledgement. The 95% upper limit is given for the single-harmonic search along with the mass quadrupole and ellipticity upper limits calculated using the distance listed in the table and a fiducial moment of inertia kg m2. Uncertainties on these parameters are not taken into account, and for reference, we report the used best-fit distance values in Table Acknowledgement provided by the EM observation. However, for pulsars that did not surpass their spin-down limits, these and upper limits are unphysical since they would lead to spin-down rates that are greater than their measured values. From the upper limit on the amplitude, we also compute the spin-down ratio as with defined in Equation (6). The upper limits for the dual-harmonic search are included as and . Finally, for the Bayesian method, the odds of a coherent signal versus incoherent noise are given for both the single and dual-harmonic searches. For the -statistic and for the 5-vector method, to assess the statistical significance of a candidate and quantify the consistency with the assumption of just noise, we report the p-value.
@@
As shown in Table Acknowledgement, there is broad agreement among the different pipelines, despite these pipelines being largely independent, and the statistical procedures used to derive the upper limits are different. The data frameworks and pre-processing procedures used by each pipeline account for the differences found in the upper limit results.
@@ 6 Discussion
In this section, we discuss the results in Table Acknowledgement. Motivated by the comparable results among the three pipelines, we consider the Bayesian pipeline as a reference. As described in the previous section, we have no evidence of a CW signal in any of the searches we conducted.
@@
34 pulsars out of the 45 considered targets in Table Acknowledgement have been already analyzed in the joint O2 plus O3 analysis (Abbott et al., 2022). Overall the corresponding upper limits on the GW amplitude are comparable for the single-harmonic search, with some targets showing better results in O4a and some targets having worse results than those in Abbott et al. (2022), see Figure 6. This is expected since the targeted search sensitivity of the O2–O3 dataset is comparable to the O4a one, except at very low frequencies. The targeted search sensitivity can be expressed in terms of minimum detectable amplitude (D’Onofrio et al., 2024), . For a multi-detector analysis considering detectors and averaging over the sky position and polarization parameters,
where the factor (the exact value depending on the considered pipeline), while and are respectively the effective observation time and the average power spectral density (PSD) for the -th detector. For the O4a targeted search sensitivity (with an observation time of approximately seconds for both detectors), see the pink curve in Figure 2.
@@
Theoretical estimates of the maximum mountain sizes that an elastically deformed neutron star can sustain are subject to significant uncertainties, with estimates for the ellipticity ranging from for conventional neutron stars, to as large as for stars with exotic solid phases (see e.g. Ushomirsky et al. (2000); Owen (2005); Haskell et al. (2007); Johnson-McDaniel & Owen (2013); Gittins & Andersson (2021); Morales & Horowitz (2022)). Comparison with the results given in Figure 4 and Table Acknowledgement show that our observationally-obtained upper limits overlap with these ranges, confirming we are continuing to push into the regime of astrophysical interest. Estimates for magnetically-induced ellipticities are similarly uncertain (see e.g. Haskell et al. (2008); Glampedakis et al. (2012); Fujisawa et al. (2022)). Nevertheless, to give a concrete example, Dall’Osso & Perna (2017) (see however Lander & Jones (2018)) have suggested that the apparent gradual increase in the angle between the spin axis and magnetic axis of the Crab pulsar provides evidence for a magnetically-induced ellipticity . This to be compared with our upper limit of for the Crab, again confirming we are probing a regime for astrophysical interest.
@@ Acknowledgement
Note. — For references and other notes see Table Acknowledgement. Values in parentheses are those produced using the restricted orientation priors described in Section 3.1.1. The last column shows the false-alarm probability (FAP) for a signal, assuming that the value has a distribution with 2 degrees-of-freedom.

### v2

@@ 4.2 EM dataset
For this analysis, we selected pulsars with a rotation frequency close to or greater than , to lie in the bandwidth of the LIGO detectors, and with an expected targeted search sensitivity for the strain amplitude within a factor 3 of the spin-down limit (see Figure 1 for the pulsars’ sky locations and Table 2 for the list of analyzed pulsars). Out of the 45 pulsars analyzed in this work, 11 pulsars belong to binary systems and there are 10 millisecond pulsars with frequencies higher than .
@@ 5.1 Targeted searches
The results are shown in Table 2. The 95% upper limit is given for the single-harmonic search along with the mass quadrupole and ellipticity upper limits calculated using the distance listed in the table and a fiducial moment of inertia kg m2. Uncertainties on these parameters are not taken into account, and for reference, we report the used best-fit distance values in Table 2 provided by the EM observation. However, for pulsars that did not surpass their spin-down limits, these and upper limits are unphysical since they would lead to spin-down rates that are greater than their measured values. From the upper limit on the amplitude, we also compute the spin-down ratio as with defined in Equation (6). The upper limits for the dual-harmonic search are included as and . Finally, for the Bayesian method, the odds of a coherent signal versus incoherent noise are given for both the single and dual-harmonic searches. For the -statistic and for the 5-vector method, to assess the statistical significance of a candidate and quantify the consistency with the assumption of just noise, we report the p-value.
@@
As shown in Table 2, there is broad agreement among the different pipelines, despite these pipelines being largely independent, and the statistical procedures used to derive the upper limits are different. The data frameworks and pre-processing procedures used by each pipeline account for the differences found in the upper limit results.
@@ 6 Discussion
In this section, we discuss the results in Table 2. Motivated by the comparable results among the three pipelines, we consider the Bayesian pipeline as a reference. As described in the previous section, we have no evidence of a CW signal in any of the searches we conducted.
@@
34 pulsars out of the 45 considered targets in Table 2 have been already analyzed in the joint O2 plus O3 analysis (Abbott et al., 2022). Overall the corresponding upper limits on the GW amplitude are comparable for the single-harmonic search, with some targets showing better results in O4a and some targets having worse results than those in Abbott et al. (2022), see Figure 6. This is expected since the targeted search sensitivity of the O2–O3 dataset is comparable to the O4a one, except at very low frequencies. The targeted search sensitivity can be expressed in terms of minimum detectable amplitude (D’Onofrio et al., 2024), . For a multi-detector analysis considering detectors and averaging over the sky position and polarization parameters,
where the factor (the exact value depending on the considered pipeline), while and are respectively the effective observation time and the average power spectral density (PSD) for the -th detector. For the O4a targeted search sensitivity (with an observation time of approximately seconds for both detectors), see the pink curve in Figure 2. The O4 targeted searches have a sensitivity depth of around 500; here, is the power spectrum taking the harmonic mean of the data over time and over detectors, and the upper limit on the pulsar amplitude [Wette, Behnke+, Dreissigacker+].
@@
Theoretical estimates of the maximum mountain sizes that an elastically deformed neutron star can sustain are subject to significant uncertainties, with estimates for the ellipticity ranging from for conventional neutron stars, to as large as for stars with exotic solid phases (see e.g. Ushomirsky et al. (2000); Owen (2005); Haskell et al. (2007); Johnson-McDaniel & Owen (2013); Gittins & Andersson (2021); Morales & Horowitz (2022)). Comparison with the results given in Figure 4 and Table 2 show that our observationally-obtained upper limits overlap with these ranges, confirming we are continuing to push into the regime of astrophysical interest. Estimates for magnetically-induced ellipticities are similarly uncertain (see e.g. Haskell et al. (2008); Glampedakis et al. (2012); Fujisawa et al. (2022)). Nevertheless, to give a concrete example, Dall’Osso & Perna (2017) (see however Lander & Jones (2018)) have suggested that the apparent gradual increase in the angle between the spin axis and magnetic axis of the Crab pulsar provides evidence for a magnetically-induced ellipticity . This to be compared with our upper limit of for the Crab, again confirming we are probing a regime for astrophysical interest.
@@ Acknowledgement
Note. — For references and other notes see Table 2. Values in parentheses are those produced using the restricted orientation priors described in Section 3.1.1. The last column shows the false-alarm probability (FAP) for a signal, assuming that the value has a distribution with 2 degrees-of-freedom.